OpenStudy (anonymous):

Factor completely. 3x3 – 12x A.3x(x – 2)(x – 2) B.3x(x + 2)(x – 2) C.3(x + 2)(x – 2) D.3x(x – 4)

OpenStudy (solomonzelman):

3x³ - 12x, like this?

OpenStudy (anonymous):

yes

OpenStudy (decentnabeel):

\[\mathrm{Factor}\:3x^3-12x:\quad 3x\left(x+2\right)\left(x-2\right)\]

OpenStudy (solomonzelman):

ok, I will tell you that: 3x³ = 3x • x² 12x = 3x • 4

OpenStudy (decentnabeel):

answer is option B

OpenStudy (anonymous):

tanks bunches

OpenStudy (anonymous):

thanks

OpenStudy (decentnabeel):

\[\mathrm{Factor\:out\:}3x\] \[=3x\left(x^2-4\right)\] \[\mathrm{Factor}\:x^2-4:\quad \left(x+2\right)\left(x-2\right)\] \[=3x\left(x+2\right)\left(x-2\right)\]

OpenStudy (decentnabeel):

are you ok @SolomonZelman

OpenStudy (solomonzelman):

I know why I am here, to help people and not to give them answer. They maybe don't realize, but they won't have openstudy for their non-online tests. Also, I don't really see a single reason for why I shouldn't be ok.

OpenStudy (anonymous):

take a chill pill

OpenStudy (solomonzelman):

Just imagine yourself totally unprepared for an exam..... there is a way to prevent that, you know.

OpenStudy (solomonzelman):

I won't further overwhelm on you though, I don't see how my talking right now can be any useful.... good luck, do well!